OpenStudy (anonymous):

solve this equation 1/4x-3/4 = 7/x

OpenStudy (anonymous):

Multiply everything by x. It will turn it into a quadratic equation that you can solve. Make sure to check your solutions by substitution into the original equation.

OpenStudy (anonymous):

(1-3x)/4x = 7/x 1-3x = 28 3x = -27 x = -9

OpenStudy (anonymous):

saruz answered the question correctly for how it was posed in the question. I answered assuming the question is actually\[\frac{1}{4}x-\frac{3}{4} = \frac{7}{x}\]

OpenStudy (anonymous):

what the question is in your perception @AnimalAin

OpenStudy (anonymous):

I've seen so much imprecise notation here that I sort of read between the lines to figure out what the intended question is. Not sure in this case.

OpenStudy (anonymous):

Oh, Well thanks to both of you.